If 10 people commit 23 mistakes in 15 months, how many mistakes would be committed by 170 people in 12 months?
Answer by stanbon(75887)

If 10 people commit 23 mistakes in 15 months, how many mistakes would be committed by 170 people in 12 months?
----
# of mistakes(m) is directly proportional to # of people(p) and # of months(t).
--------
m = k*(pt)
---
solve for "k" using "10 people commit 23 mistakes in 14 months":
23 = k*10*14
k = 23/140
----
Equation:
m = (23/140)(pt)
----------------------
How many mistakes would be committed by 170 people in 12 months?
m = (23/140)(170*12)
m = 335 +
====================
Cheers,
Stan H.